OpenStudy (anonymous):

help cant figure out this problem 6(a+3)=21-3-(1-2a)

OpenStudy (anonymous):

Are you trying to find a

OpenStudy (jhannybean):

First apply the distributive property.

OpenStudy (anonymous):

Haha Nevermind XD

OpenStudy (jhannybean):

6*a +6*3 = 21 -3 -1*1 -1 *-2a

OpenStudy (anonymous):

i do that but the thing i get stuck on is putting it all togethter

OpenStudy (jhannybean):

Oh okay. so when we simplify this we get 6a +18 = 21 -3-1+2a right?

OpenStudy (jhannybean):

We'll work through it step by step.

OpenStudy (jhannybean):

Are you following? :\

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

instead of 2a wouldnt it be -6a

OpenStudy (jhannybean):

Let's see. 6a +18 = 21 -3-1+2a simplify the right side 6a + 18 = 21-4 +2a simplify again 6a + 18 = 17 + 2a bring all the a's to one side of the equation 6a -2a +18 = 17 simplify left side 4a +18 = 17 bring all the numbers to the other side of the equation 4a = 17 - 18 simplify 4a = -1 isolate a on one side by dividing both sides of the equation by 4 4a/4 = -1/4 Simplify a = -1/4

OpenStudy (anonymous):

okay thanks